### v1

@@ Abstract
In this work, we propose the use of Ground Penetrating Radar (GPR) for rover localization on Mars. Precise pose estimation is an important task for mobile robots exploring planetary surfaces, as they operate in GPS-denied environments. Although visual odometry provides accurate localization, it is computationally expensive and can fail in dim or high-contrast lighting. Wheel encoders can also provide odometry estimation, but are prone to slipping on the sandy terrain encountered on Mars. Although traditionally a scientific surveying sensor, GPR has been used on Earth for terrain classification and localization through subsurface feature matching. The Perseverance rover and the upcoming ExoMars rover have GPR sensors already equipped to aid in the search of water and mineral resources. We propose to leverage GPR to aid in Mars rover localization. Specifically, we develop a novel GPR-based deep learning model that predicts 1D relative pose translation. We fuse our GPR pose prediction method with inertial and wheel encoder data in a filtering framework to output rover localization. We perform experiments in a Mars analog environment and demonstrate that our GPR-based displacement predictions both outperform wheel encoders and improve multi-modal filtering estimates in high-slip environments. Lastly, we present the first dataset aimed at GPR-based localization in Mars analog environments, which will be made publicly available upon publication.
@@ 1 INTRODUCTION
Precisely localizing planetary rovers in GPS-denied environments such as the Moon and Mars is a challenging task. Until recently, localization for the Perseverance rover on the surface of Mars has required a human-in-the-loop for manual corrections and tuning [1, 2]. New approaches have seen increasingly autonomous solutions for global state estimation [3], but this problem has not been entirely solved just yet. As interest increases in launching lower-cost rovers with shorter lifespans, there is a need for real-time GPS-denied rover localization that does not require a large team of human operators in the loop.
@@
Table 1: Localizing Ground Penetrating Radar Datasets
@@
Spirit rover [6], making encoders an unreliable sensor for localization. Visual odometry (VO) has shown the benefits of including additional sensing modalities to improve state estimation. In VO, features are tracked between subsequent pairs of images in order to characterize the robot’s motion and look for loop closure. Although VO typically outperforms inertial and wheel encoder odometry, this method is more computationally costly. In a VO demonstration for Martian rovers, Spirit and Opportunity needed three minutes to complete a single vision-based tracking step [7].
Furthermore, vision-based odometry can also suffer from failure modes surrounding lighting [8] – too much light contrast can make feature tracking impossible, and performing VO in the dark is an all but impossible task.
In the space domain, scientific sensors can be creatively repurposed to provide information about a robot’s environment to the navigation algorithms, such as terrain type or slip characterization [9, 10]. Ground Penetrating Radar (GPR) has already demonstrated promise on Earth as a tool for localization by using subsurface features [11, 12] and for terrain classification using deep learning methods [13]. The Perseverance rover is equipped with the RIMFAX [14] GPR sensor, and the planned European Space Agency ExoMars rover will have a GPR to search for in-situ drilling sites [15].
@@ 2.2 Ground Penetrating Radar
GPR is a popular sensor in the fields of civil engineering [19], archaeology [20], environmental science [21], and planetary geology [22]. The GPR on-board the Perseverance rover has been used to search for water and to characterize the electromagnetic properties of the Martian subsurface [14]. The signal returns from a GPR allow for a nonintrusive view of subsurface materials and features. Single GPR returns, or a-scans, can be concatenated sequentially to form a radargram image (or b-scan) that encodes 2D spatial information.
Subsurface features are generally more static than visual features above the surface – for example, a scene might look completely different to a camera after a fresh coat of snow, or during a rainstorm. Additionally, processing GPR traces is less computationally intensive than performing visual feature matching. Relying on GPR features as an additional mode of localization could improve absolute positioning of planetary rovers.
@@ 2.3 Localizing Ground Penetrating Radar
The use of GPR for localization has been under development for the last decade. Although some work has investigated the use of matching single GPR a-scans [23], most approaches take advantage of patterns in sequential GPR traces that are present in concatenated b-scans. Early approaches rely on a GPS prior and a pre-registered map in order to match incoming GPR measurements to an existing database of underground features [24, 25]. In order to reduce reliance on GPS, subsequent methods [11, 26, 27, 28, 29, 30, 31, 32] focused on more robust approaches for GPR feature detection that could aid in matching newly collected data with the pre-registered map of GPR traces. Some approaches utilized signal processing for finding and matching features across b-scans [11, 30, 31, 32], while other works use deep learning for feature extraction [26, 27, 28, 29].
@@ 3.1 Ground Penetrating Radar Sensor Model
GPR uses radio waves to characterize subsurface properties, often depicted as a time series radargram (Figure 2). Although a radargram might appear to be a simple pictoral cross-section of the area below the sensor, the sensor returns do not directly capture a visual “snapshot” of subterranean feature geometry. This is due to fact that GPR samples capture relative permittivity between materials [34].
@@
Figure 2: Consecutive GPR samples have overlapping beam spreads, which cause them to detect the same object across multiple returns. Because of this, the resulting GPR radargram image may not intuitively represent the geometry of objects underground. However, we can use this property of the sensor to predict relative pose translation using deep learning.
Figure 3: Our GPRFormer network takes in a set of $k$ consecutive GPR a-scans as vectors, encodes them into a latent space, and then passes the encoded tokens through a compact transformer [35]. The output from the network is an estimate of relative displacement ($\Delta d$) that the vehicle traveled while collecting the GPR traces in the input.
However, the shape of the object visible in the radargram may not intuitively match the true geometry of the object. A GPR has a wide beam dispersion, so a submerged object will reflect some of the signal back at the transmitter even if the GPR is not directly over the object as shown in Figure 2. This results in an almost imperceptable correlation of consecutive a-scans. It is this idea that has inspired our investigation into using deep learning for predicting relative pose translation from GPR returns.
@@
In order to reduce signal noise, GPR sensors often collect several samples immediately after one another and average them in a process called “stacking.” The stacking process is usually applied during signal postprocessing [34]. Another method to reduce signal noise and emphasize visible features is the use of signal filtering, including dewow filters, triangular bandpass filters, spreading and exponential compensation, and wavelet denoising [5].
@@ 3.2 GPRFormer
Inspired by the body of work looking to solve 3D relative camera pose estimation [37], we use a supervised compact transformer with multi-head cross attention to detect similarities between a batch of $s$ sequential GPR a-scans. The encoder ensures that our input raw data is transformed to a meaningful latent space representation for the transformer to encode further. Within the compact transformer, multi-head attention is excellent at encoding nuanced relationships between tokens, which in our case are individual GPR a-scans. Finally, a regressor predicts the final displacement through a fully connected layer. A Mean Squared Error (MSE) loss is used during training.
@@ 3.3 Extended Kalman Filter for Localization
With an assumption that the robot operates in locally 2D environments in the East North Up (ENU) frame, we denote the robot’s state ($\mathbf{x}_{t}$) and measurement ($\mathbf{z}_{t}$) vectors as:
@@
[…]
@@
To predict the current state given the previous state, the nonlinear model is used. We can also propagate the covariance, $\boldsymbol{P}$, using the Jacobian of the nonlinear model $\boldsymbol{F}$ and the process noise covariance $\boldsymbol{Q}$:
$\hat{\mathbf{x}}_{k}$ $=f(\mathbf{x}_{k-1})$ (5)
$\hat{\boldsymbol{P}}_{k}$ $=\boldsymbol{F}\boldsymbol{P}_{k-1}\boldsymbol{F}^{\top}+\boldsymbol{Q}.$ (6)
The second step, an update given the measurement corrections, is calculated using $\hat{\mathbf{x}}_{k}$, $\hat{\boldsymbol{P}}_{k}$, the observation matrix $\boldsymbol{H}$, Kalman gain $\boldsymbol{K}$, and the measurement covariance $\boldsymbol{R}$.
$\boldsymbol{K}$ $=\hat{\boldsymbol{P}}_{k}\boldsymbol{H}^{\top}\left(\boldsymbol{H}\hat{\boldsymbol{P}}_{k}\boldsymbol{H}^{\top}+\boldsymbol{R}\right)^{-1}$ (7)
$\mathbf{x}_{k}$ $=\hat{\mathbf{x}}_{k}+\boldsymbol{K}\left(\mathbf{z}-\boldsymbol{H}\hat{\mathbf{x}}_{k}\right)$ (8)
$\boldsymbol{P}_{k}$ $=\left(\boldsymbol{I}-\boldsymbol{K}\boldsymbol{H}\right)\hat{\boldsymbol{P}}_{k}\left(\boldsymbol{I}-\boldsymbol{K}\boldsymbol{H}\right)^{\top}+\boldsymbol{K}\boldsymbol{R}\boldsymbol{K}^{\top}$ (9)
@@ 3.3.3 GPR Relative Displacement
Figure 5: Visual comparison of geologic features between Mars and the MDRS region.
Figure 6: Areas of dataset collection in the region surrounding MDRS in southeast Utah. Note the variety of terrain types, which provide a wide distribution of GPR signals for training the relative displacement network. A snapshot sample of a GPR radargram at each site is shown in the terrain image cutout. Base MDRS map provided by the Mars Society, and cutaway map of the United States provided by OpenStreetMap.
@@ 4 DATASET
Table 2: The cumulative displacement from the ground truth and the GPRFormer network predictions using the RMSE metric in meters on the CMU-GPR dataset. Note that the CMU-GPR dataset was collected on flat concrete surfaces with little slip, making the wheel encoder very accurate. Bold is best.
Method Sequence 1 Sequence 3 Sequence 7 Sequence 9 Sequence 12-17-01-44 Sequence 15-14-58-25
Wheel Encoder 0.477 0.412 0.364 0.130 1.840 0.930
GPRFormer (Ours) 2.895 3.661 5.9945 5.182 0.915 0.668
@@
Note that we also collected GPS data on a Garmin 18x unit, however, due to an issue in parsing of satellite-based augmentation signals, the GPS data is not good enough to provide ground truth localization. Instead, we provide localizations from a map built and optimized offline by RTAB-Map [39], a popular visual odometry and 3D reconstruction tool. Visual odometry is a well-posed task for the images in our dataset, as there is little camera motion blur and a large number of trackable features. These reference poses from VO are optimized offline, which is a computationally costly task and is infeasible to run in real-time on spaceflight hardware. However, we are able to use it as a good reference baseline for our experiments in the absence of GPS.
In order to prepare the MarsLGPR dataset for training, we first remove anomalous traces where the recorded values are outside of the ±$\pm 50$ mV range and then we average the consecutive stacked traces to help reduce noise. Additionally, we perform linear and SLERP interpolation along the RTAB-Map poses to have time synchronized sensor and positional data, which is important for model training.
@@ 5 EXPERIMENTS & RESULTS
In this section, we present experiments to validate our method qualitatively and quantitatively. We first present accuracy metrics and ablations on GPRFormer, our novel GPR-based relative displacement network. Then, we qualitatively and quantitatively evaluate the GPRFormer + EKF localization framework. Experiments are performed on both the publicly available CMU-GPR dataset [5] and on our own MarsLGPR dataset.
Figure 8: Comparison of GPRFormer relative displacement predictions against the wheel encoders and the ground truth on four test sequences from the CMU-GPR [5] The left column shows the cumulative displacement over time, and the right column shows the accompanying predictions for the displacement at each time step.
Figure 9: Comparison of GPRFormer relative displacement predictions against the wheel encoders and RTAB-Map on two test sequences from the MarsLGPR dataset. The left column shows the cumulative displacement over time, and the right column shows the accompanying predictions for the displacement at each time step. Note that the wheel encoders accumulate more displacement over time. This is due to the slippery terrain causing the wheels to spin out.
5.1 Implementation Details
We train two versions of our GPRFormer model, one on the CMU-GPR dataset and one on the MarsLGPR dataset. For the CMU-GPR dataset, we follow [33] and use sequences 0, 2, 4, 5, 6, 8, and 10 for training, and sequences 1, 3, 7, and 9 for testing. We apply background removal, dewow filtering, spreading and exponential compensation, and wavelet denoising to enhance feature visibility. During training, the model is supervised with displacement calculated from the ground truth RTK GPS positioning data for the CMU-GPR dataset and from reference RTAB-Map poses for the MarsLGPR dataset.
The Adam optimizer is used, and a linear learning rate scheduler helps with better fine-tuning later in the training process. All training is performed on an NVIDIA GeForce 32 GB RTX 3090 GPU and takes around 10 minutes to complete for each dataset.

### v2

@@ Abstract
In this work, we propose the use of Ground Penetrating Radar (GPR) for rover localization on Mars. Precise pose estimation is an important task for mobile robots exploring planetary surfaces, as they operate in GPS-denied environments. Although visual odometry provides accurate localization, it is computationally expensive and can fail in dim or high-contrast lighting. Wheel encoders can also provide odometry estimation, but are prone to slipping on the sandy terrain encountered on Mars. Although traditionally a scientific surveying sensor, GPR has been used on Earth for terrain classification and localization through subsurface feature matching. The Perseverance rover and the upcoming ExoMars rover have GPR sensors already equipped to aid in the search of water and mineral resources. We propose to leverage GPR to aid in Mars rover localization. Specifically, we develop a novel GPR-based deep learning model that predicts 1D relative pose translation. We fuse our GPR pose prediction method with inertial and wheel encoder data in a filtering framework to output rover localization. We perform experiments in a Mars analog environment and demonstrate that our GPR-based displacement predictions both outperform wheel encoders and improve multi-modal filtering estimates in high-slip environments. Lastly, we present the first dataset aimed at GPR-based localization in Mars analog environments, which will be made publicly available at https://umfieldrobotics.github.io/marslgpr/.
@@ 1 INTRODUCTION
Precisely localizing planetary rovers in GPS-denied environments such as the Moon and Mars is a challenging task. Until recently, localization for the Perseverance rover on the surface of Mars has required a human-in-the-loop for manual corrections and tuning [1, 2]. New approaches have seen increasingly autonomous solutions for global state estimation [3], but this problem has not been entirely solved. As interest increases in launching lower-cost rovers with shorter lifespans, there is a need for real-time GPS-denied rover localization that does not require a large team of human operators in the loop.
@@
Table 1: Localizing Ground Penetrating Radar datasets.
@@
Spirit rover [6], making encoders an unreliable sensor for localization. Visual odometry (VO) has shown the benefits of including additional sensing modalities to improve state estimation. In VO, features are tracked between subsequent pairs of images in order to characterize the robot’s motion and look for loop closure. Although VO typically outperforms inertial and wheel encoder odometry, this method is more computationally costly and is not realistic for rovers to run in real-time given current hardware constraints. In a VO demonstration for Martian rovers, Spirit and Opportunity needed three minutes to complete a single vision-based tracking step [7]. The Curiosity rover has demonstrated on-board VO tests, but with each image processing step taking around 34 seconds [8]. For reference, state-of-the-art terrestrial VO runs at 30-40 frames per second [9]. The Perseverance rover localization framework [3] includes offline optimization-based localization methods, such as matching against existing orbital elevation maps, but does not mention use of real-time VO on board the rover.
In the space domain, scientific sensors can be creatively repurposed to provide information about a robot’s environment to the navigation algorithms, such as terrain type or slip characterization [10, 11]. Ground Penetrating Radar (GPR) has already demonstrated promise on Earth as a tool for localization by using subsurface features [12, 13] and for terrain classification using deep learning methods [14]. The Perseverance rover is equipped with the RIMFAX [15] GPR sensor, and the planned European Space Agency ExoMars rover will have a GPR to search for in-situ drilling sites [16]. The upcoming CADRE mission to the surface of the Moon plans to have several robots – each equipped with GPR – collect data cooperatively [17].
@@ 2.2 Ground Penetrating Radar
GPR is a popular sensor in the fields of civil engineering [21], archaeology [22], environmental science [23], and planetary geology [24]. The GPR on-board the Perseverance rover has been used to search for water and to characterize the electromagnetic properties of the Martian subsurface [15]. The signal returns from a GPR allow for a nonintrusive view of subsurface materials and features. Single GPR returns, or A-scans, can be concatenated sequentially to form a B-scan image that encodes 2D spatial information.
Subsurface features are generally more static than visual features above the surface – for example, a scene might look completely different to a camera after a fresh coat of snow or during a rainstorm. Additionally, processing GPR traces is less computationally intensive than performing visual feature matching. Relying on GPR features as an additional mode of localization could improve absolute positioning of planetary rovers.
@@ 2.3 Localizing Ground Penetrating Radar
The use of GPR for localization has been under development for the last decade. Although some work has investigated the use of matching single GPR A-scans [25], most approaches take advantage of patterns in sequential GPR traces that are present in concatenated B-scans. Early approaches rely on a GPS prior and a pre-registered map in order to match incoming GPR measurements to an existing database of underground features [26, 27]. In order to reduce reliance on GPS, subsequent methods [12, 28, 4, 29, 30, 31, 32, 33] focus on more robust approaches for GPR feature detection that could aid in matching newly collected data with the pre-registered map of GPR traces. Some approaches utilize signal processing for finding and matching features across B-scans [12, 31, 32, 33], while other works use deep learning for feature extraction [28, 4, 29, 30].
@@ 3.1 Ground Penetrating Radar Sensor Model
GPR uses radio waves to characterize subsurface properties, often depicted as a time series B-scan (Figure 2). Although a B-scan might appear to be a simple pictoral cross-section of the area below the sensor, the sensor returns do not directly capture a visual “snapshot” of subterranean feature geometry. This is due to the fact that GPR samples capture relative permittivity between materials [35].
@@
Figure 2: Consecutive GPR samples have overlapping beam spreads, which cause them to detect the same object across multiple returns. Because of this, the resulting GPR B-scan image may not intuitively represent the geometry of objects underground. However, we can use this property of the sensor to predict relative pose translation using deep learning.
Figure 3: Our GPRFormer network takes in a GPR B-scan with $k$ consecutive traces, performs a linear encoding, adds learned positional embeddings, and then passes the encoded tokens through a compact transformer [36]. The output from the network is an estimate of relative displacement ($\Delta d$) that the vehicle traveled while collecting the GPR traces in the input.
However, the shape of the object visible in the B-scan may not intuitively match the true geometry of the object. A GPR has a wide beam dispersion, so a submerged object will reflect some of the signal back at the transmitter even if the GPR is not directly over the object as shown in Figure 2. This results in an almost imperceptible correlation of consecutive A-scans. It is this idea that has inspired our investigation into using deep learning for predicting relative pose translation from GPR returns.
@@
3.1.1 Filtering
In order to reduce signal noise and emphasize visible features, signal filtering such as background removal, dewow filters, Spreading and Exponential Compensation (SEC) gain, and wavelet denoising may be utilized [5]. An example of pre- and post-filtered B-scan is shown in Figure 4.
Background removal is one of the most standard GPR de-noising approaches. It involves removing “background” horizontal layering artifacts, typically produced from a direct wave [35]. We use a global averaging window over the entire batch of traces $\mathbf{G}=[\mathbf{g}_{0},\mathbf{g}_{1},\dots,\mathbf{g}_{k-1}]$ where $k$ is the number of traces in the network input:
$\mathbf{G}_{ij}^{\prime}=\mathbf{G}_{ij}-\frac{1}{k}\sum_{j=0}^{k-1}\mathbf{G}_{ij}$ (2)
Dewow filters remove low-frequency drift from GPR data with polynomial detrending [37]. For a given trace $\mathbf{g}\in\mathbb{R}^{t}$, we find the coefficients $\mathbf{a}$ of a degree-$\rho$ polynomial representing the drift for each trace $\mathbf{g}\in\mathbf{G}$, where $\rho=3$.
$\mathbf{a}=\arg\min\sum_{i=1}^{t}\left|\;\mathbf{g}_{i}-\sum_{j=0}^{p}a_{j}i^{j}\;\right|^{2}$ (3)
$\mathbf{p}_{i}=\sum_{j=0}^{\rho}a_{j}i^{j},\;\;\text{ for }i=0,1,\dots,t-1$ (4)
$\mathbf{g}^{\prime}=\mathbf{g}-\mathbf{p}$ (5)
Figure 4: A comparison of a raw B-scan containing $k=10$ traces (a.) and the filtered result after applying background removal, dewow filtering, SEC gain, and wavelet denoising (b.). Note that the direct wave bands are removed and the deeper features (more rightwards) are more pronounced.
SEC gain compensates for the loss of energy as the wavefront propagates through the ground. It does this by scaling the amplitude depending on the order of return [38]. Parameter $a$ controls the magnitude of the amplification, while parameter $b$ controls how exponential or linear the amplification is. Threshold $T_{\gamma}$ sets a maximum amplification at a certain index to avoid over-emphasizing deeper returns (especially when $b$ is closer to 0, and is thus more exponential). We set $a=0.015$, $b=0$, and $T_{\gamma}=100$.
$\boldsymbol{\gamma}=\begin{cases}i^{b}\,e^{a\,i}&\text{if }i<T_{\gamma}\\ {T_{\gamma}}^{b}\,e^{aT_{\gamma}}&\text{if }i\geq T_{\gamma}\\ \end{cases}\;\;\text{ for }i=0,1,\dots,t-1$ (6)
$\boldsymbol{\gamma}$ is then multiplied, element-wise, across each row in the B-scan $\mathbf{G}$ [5].
Wavelet denoising aims to separate out the true GPR signal from noise by decomposing the input into the wavelet domain, applying thresholding, and then reconstructing the input in the image space. We selected the Daubechies-6 “mother” wavelet due to its similarity to a GPR signal and previously-demonstrated superior performance [39].
In order to reduce signal noise, GPR sensors often collect several samples immediately after one another and average them in a process called “stacking.” The stacking process is usually applied during signal postprocessing [35]. We apply stacking over sets of three traces.
@@ 3.2 GPRFormer
Inspired by methods for 3D relative camera pose estimation [41], we use a supervised compact transformer architecture to process temporal sequences of 1D sensor data, concatenated into a GPR B-scan. The model consists of three primary components: a linear input encoder that maps 200-dimensional raw sensor readings to 256-dimensional token embeddings, a modified CCT backbone with dual sequence pooling, and a regression head for displacement prediction. Mean Squared Error (MSE) loss is used during training.
3.2.1 Linear Input Encoder
The input encoder is a simple linear projection layer that is applied to each of the $k$ consecutive sensor readings to transform them from raw 200-dimensional feature space to a 256-dimensional embedding space. The GPR B-scan format is natively structured and very compact, so we found this simple encoder architecture to be sufficient.
A key component of the model is the positional embedding scheme, which allows the transformer to understand the sequential nature of the input data, as the traces are stacked in order of capture. We specifically utilize learned positional embeddings, which allows the model to adapt positional representations specifically for the sensor data domain. The positional embeddings are added element-wise to the token embeddings before entering the transformer stack.
3.2.2 CCT Transformer
Our modified CCT backbone has six transformer encoder layers, each containing pre-normalization, multi-head self-attention with four attention heads, and feed-forward networks each with an expansion ratio of 3.0. The architecture includes a dual sequence pooling mechanism that combines features from both pre-transformer and post-transformer representations. Two separate attention pooling layers generate sequence-wise attention weights, producing pooled representations $\mathbf{x}_{1}$ (from post-transformer features) and $\mathbf{x}_{2}$ (from pre-transformer features applied to post-transformer outputs). The final representation combines these via a learnable parameter $\alpha$:
$x_{\text{final}}=\alpha\mathbf{x}_{1}+(1-\alpha)\mathbf{x}_{2}.$ (7)
This allows the model to balance between raw sequential patterns and learned contextual representations from the transformer.
3.2.3 Regression Head
The original CCT classification head is replaced with an identity layer, and finally a regression head consisting of two fully connected layers with ReLU activation and dropout ($p=0.1$) predicts the final single displacement distance.
@@ 3.3 Extended Kalman Filter for Localization
We make a simplifying assumption that the robot is operating in locally 2D environments in the East North Up (ENU) frame. The sensor mounted on the rear of the robot platform limits its movements to relatively planar regions. We denote the robot’s state ($\mathbf{x}_{t}$) and measurement ($\mathbf{z}_{t}$) vectors as:
@@
where $t$ is the continuous time, $(x,y)$ is the position, $\psi$ is the yaw angle, $\cdot$ denotes velocity and $\cdot\cdot$ denotes acceleration. The estimation of a robot’s full pose at discrete timestep $k$ can be described with the state equation:
@@ 3.3.3 GPR Relative Displacement
Figure 6: Visual comparison of geologic features between the MDRS region (top) and Mars (bottom).
Figure 7: Areas of dataset collection in the region surrounding MDRS in southeast Utah. Note the variety of terrain types, shown in the red cut-outs, which provide a wide distribution of GPR signals for training the relative displacement network. A snapshot sample of a filtered GPR B-scan at each site is shown in the terrain image cutout. Base MDRS map provided by the Mars Society, and cutaway map of the United States provided by OpenStreetMap.
@@ 4 DATASET
Table 2: The cumulative displacement from the RTAB-Map (reference), Wick.*, wheel encoder, and GPRFormer network predictions using the RMSE metric in millimeters on the CMU-GPR dataset. Note that the CMU-GPR dataset was collected on flat concrete surfaces with little slip, making the wheel encoder very accurate. Bold is best.
Method Seq 1 Seq 3 Seq 7 Seq 9 12-17-01-44 15-14-58-25 15-15-42-09 15-15-55-14
Wheel Encoder 9.72 4.12 4.20 3.55 20.65 27.42 29.89 29.62
Wick.* [34] 22.02 18.32 14.54 14.09 - 23.56 22.46 19.37
GPRFormer (Ours) 8.62 7.64 6.51 5.98 15.59 13.32 16.92 17.58
Note that we also collected GPS data on a Garmin 18x unit, however, due to an issue in parsing of satellite-based augmentation signals, the GPS data is not good enough to provide ground truth localization. Instead, we provide localizations from a map built and optimized offline by RTAB-Map [43], a popular VO and 3D reconstruction tool. VO is a well-posed task for the images in our dataset, as there is little camera motion blur and a large number of trackable features. These reference poses from VO are optimized offline, which is a computationally costly task and is infeasible to run in real-time on spaceflight hardware. However, we are able to use it as a reference trajectory for our model training and experiments in the absence of GPS.
In order to prepare the MarsLGPR dataset for training, we first remove anomalous traces where the recorded values are outside of the $\pm 50$ mV range and then we average the consecutive stacked traces to help reduce noise. Additionally, we perform linear and SLERP interpolation along the RTAB-Map poses to have time synchronized sensor and positional data, which is important for model training.
@@ 5 EXPERIMENTS & RESULTS
In this section, we present experiments to validate our method qualitatively and quantitatively. We first present accuracy metrics on GPRFormer and two baselines: the wheel encoders and localization network Wickramanyake [34]. Then, we qualitatively and quantitatively evaluate the GPRFormer + EKF localization framework. Additionally, we include an ablation study and runtime metrics on GPRFormer. Experiments are performed on both the publicly available CMU-GPR dataset [5] and on our own MarsLGPR dataset.
Figure 9: Comparison of GPRFormer relative displacement predictions against Wick.* [34], the wheel encoders, and the ground truth on four test sequences from the CMU-GPR [5] The left column shows the cumulative displacement over time, and the right column shows the accompanying predictions for the displacement at each time step.
Figure 10: Comparison of GPRFormer relative displacement predictions against the wheel encoders Wick.* baseline, and RTAB-Map on four test sequences from the MarsLGPR dataset. The left column shows the cumulative displacement over time, and the right column shows the accompanying predictions for the displacement at each time step. Note that the wheel encoders accumulate more displacement over time. This is due to the slippery terrain causing the wheels to spin out.
5.1 GPRFormer Implementation Details
We train two versions of our GPRFormer model, one on the CMU-GPR dataset and one on the MarsLGPR dataset. For the CMU-GPR dataset, we follow [34] and use sequences 0, 2, 4, 5, 6, 8, and 10 for training, and sequences 1, 3, 7, and 9 for testing. We apply background removal, dewow filtering, spreading and exponential compensation, and wavelet denoising to enhance feature visibility. During training, the model is supervised with displacement calculated from the ground truth RTK GPS positioning data for the CMU-GPR dataset and from reference RTAB-Map poses for the MarsLGPR dataset. The MarsLGPR train set contains 35 sequences and the test set contains four sequences.
The Adam optimizer is used with a batch size of 128, and a linear learning rate scheduler helps with better fine-tuning later in the training process. All training is performed on an NVIDIA GeForce 32 GB RTX 3090 GPU.